So my question is how I would solve this proportion:
2x/4=5x+6/3
I also wanted to check if I did this proportion correctly:
5/3x=7
I multiplied the 5/3 by 3/5 and then I got an answer of x=7. Was I supposed to multiply 5/3x by 1/5? Thank you for helping!

Answer by Boreal(15235)

2x/4=(5x+6)/3
Cross multiply and get 6x=20x+24
-14x=24
x=-12/7
========
-24/28=(-60+42)/7*3
-6/7=-18/21 checks